OpenStudy (anonymous):

find the hcf of 867 and 225

Parth (parthkohli):

List out the factors and then multiply the common factors.

OpenStudy (anonymous):

ya but not getting

Parth (parthkohli):

225 = 3 * 3 * 5 * 5 867 = 3 * 17 * 17

Parth (parthkohli):

The common factor is just 3. So that's the HCF :)

OpenStudy (anonymous):

ans is????

Parth (parthkohli):

I just told you.

OpenStudy (anonymous):

thnx